OpenStudy (anonymous):

Determine Whether The Equation is always, sometimes, or never true. question 1 6(x+1)=2(5+3x) question 2 3(y+3)+5y=4(2y+1)+5

OpenStudy (anonymous):

First you need to distrubute. So let's start with the first one. 6(x+1)=2(5+3x) 6x+1=10+3x -3x -3x 3x+1=10 -1 -1 3x =9 /3 /3 x=3 So it is sometimes true because they dont equal each other but it does have a solution

OpenStudy (anonymous):

@christhebest123 do you think you understand it well enough to do the second one?

OpenStudy (anonymous):

wouldnt it be 6x+6=10+6x @Broganleab

OpenStudy (anonymous):

if we were distributing

OpenStudy (anonymous):

Oh pellet I forgot the 2. Yes that is my bad let me redo it 6(x+1)=2(5+3x) 6x+1=10+6x -6x -6x and the 6 cancels out so now you have 1=10 do you think that is always sometimes or never true?

OpenStudy (anonymous):

never true

OpenStudy (anonymous):

Yep so do the same for the second and if you fine a solution its sometimes true if there is no solution its never true and if one equals the other say like 17=17 then it is always true

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome!